If you have a 5% solution and a 50 % solution how much of each solution do you need to make 20ml of 30% solution?
Answer by Earlsdon(6103)

Let x = the number of ml of the 5% solution required, then 20-x = the number of ml of the 50% solution needed. The sum of these will equal 20 ml of 30% solution.
From these facts, we can write the necessary equation to find x, after converting the percentages to their decimal equivalents.
x(0.05)+(20-x)(0.5) = 20(0.3) Simplify and solve for x.
0.05x+10-0.5x = 6
-0.45x+10 = 6
-0.45x = -4
x = 8.89
So you would need 8.89 ml of the 5% solution and 20-x = 11.11 ml of the 50% solution to obtain 20 ml of 30% solution.